# Impact of trace amounts of water on the stability of Micro-Pattern Gaseous Detectors

### v1

@@ 1 Introduction
Our measurements are performed at the absolute humidity levels of (parts per million, proportion by volume), which correspond to a relative humidity (RH) value of up to 20% at normal temperature and pressure conditions. It should be noted that several groups reported various instabilities in MPGD detectors appearing at much larger relative humidity values while operating/testing a detector structure in an air atmosphere in the laboratory. In [11], instabilities such as increased leakage current and spark rate were observed in GEMs operated in ambient, cleanroom conditions (RH50%). Similar observations have also been made with other types of MPGDs (see, for example, [12] for resistive Micromegas or recent results with µRGroove presented in [13]). For this reason, the humidity limits for the quality control procedures of MPGDs are usually put below the RH value of 50% or even below 10%, as required in the recently concluded large-scale productions for the LHC detector upgrades [14, 15, 16]. This is well in line with the range considered in this work. In the presented studies, we focus on the fundamental stability limits of MPGD-based detectors in a gas mixture with trace amounts of water, and the regions of large RH () are not considered here.
@@ 2.1 Gas system
The gas mixture used in the presented studies is Ar-CO2 (90-10). A dedicated setup is built to study the effect of humidity on MPGDs, in which a constant amount of water vapour is reliably introduced to the gas. Figure 1 demonstrates the humidification process. The original gas mixture flows through a piping system, splitting into two separate lines. One line incorporates a bubbler filled with distilled water, allowing the gas to absorb humidity without introducing additional contaminants like oxygen. Different humidity levels can be achieved by adjusting the flow through the bubbler using a mass flow meter, as shown in Fig. 2. Simultaneously, the second line directs dry gas around the humidifying line until they merge. The resulting gas mixture is flushed through the detector chamber while a rotameter in the dry gas line regulates and monitors the flow. A Rapidox 3100 Multigas Analyser from Cambridge Sensotec is placed after the detector vessel to record the oxygen and humidity content of the mixture. The oxygen content is kept at a minimum ( ppmV) throughout all measurements. The ambient conditions are tracked with thermometers and a pressure sensor.
@@
Pitch | Density | Transparency | Producer | |||||||
@@ 2.4 Gain measurements
The absolute gain is defined as the ratio of the amplification current to the primary current measured in the drift volume. The currents are measured using a floating multi-channel picoamperemeter [19]. For the (TH)GEM, this process is carried out with zero induction field, while the amplified signal from avalanche multiplication is read at the bottom of the foil. The primary current is measured at the top (TH)GEM electrode, with the bottom electrode and the anode plane grounded. A drift field of is applied, at which the electron collection efficiency of primary charges is expected to be nearly 100% [20, 21, 22, 23]. The measurement with a Micromegas is performed similarly by applying a drift field of and keeping the mesh and the anode plane grounded. Since the electron transparency of MMG1 is known to be less than 100% for the measured field configurations, the gain values are corrected using data obtained with this specific Micromegas structure [17], where electron transparency is measured as a function of the ratio , with representing the field strength in the amplification region. For MMG3, a 100% collection efficiency can be safely assumed [17].
@@ 2.5 Amplitude spectra
The energy resolution values are calculated by fitting a Gaussian to the main peak of the spectrum and taking the ratio of the Gaussian width () to the mean peak position . The systematic uncertainty is estimated by varying the fit ranges. The mean peak position is then corrected for the temperature and pressure () variations assuming linear dependency of gain (thus value) on ratio, which was monitored throughout the measurement. It should be noted that no detailed calibration is available for the detectors under study. Thus, although the variations were rather limited during the measurement time, the residual gain dependency remains uncorrected.
@@
The GEM gain dependency is less apparent in the measurement with decreasing humidity. Apart from the residual influence, this may also be related to water absorbed in polyimide during the first measurement series, which influences the electrical and electrostatic properties of the GEM structure (e.g. charging-up behaviour). Indeed, in the case of MMG3, no significant gain dependency and no differences are observed between the measurements with increasing and decreasing humidity content. The Micromegas pillars are made of photopolymer solder mask material (DuPontTM Pyralux® PC 1000) and cover roughly of the detector area ([30]), contrary to GEMs, where polyimide foil (Apical®) fills more than 75% of the active area (see Table 1). Thus the water uptake by the MMG3 detector shall be largely reduced in comparison to GEMs, which may explain the different results obtained with these two types of detectors. In addition, the hygroscopicity of GEM and MMG3 materials shall be considered, as they can further influence the water trapping in these two structures. It is expected that the water absorption of a polymer mask is lower than that of the Apical® foil (water absorption of 2.7%, according to the producer [31]), however, no exact value could be identified for the former.
@@
For completeness, the effect of gas humidity on energy resolution is studied. It is observed to be very weak for both GEM and MMG3, even though, given the electronegativity of water molecules, one could expect the attachment to deteriorate the energy resolution. As discussed above, the influence of humidity on the effective Townsend coefficient for different gases may vary depending on the electric field strength, however, we do not consider this effect in the current studies. The GEM measurements show a slight improvement in the energy resolution with increasing water content which may be related to the increasing gain values (see above). Still, the relative change of energy resolution by less than 5% in the measured humidity range can be considered marginal. To summarise, the effect of humidity on the performance of the detector is found to be negligible and should not alter the studies on discharge stability. Of course, a complete interpretation of the observed effects should be considered with a dedicated measurement campaign.
@@ 2.6 Discharge measurement
The discharge rate is defined as a number of measured discharges occurring within the measurement time . Consequently, the discharge probability is given by the ratio of the discharge rate to the measured source rate , following:
Counting the discharges is done by installing an open cable to the detector chamber, acting as an antenna. The cable is connected to an oscilloscope, which counts and visualises the sparks as shown in figure Fig. 6.
To measure the discharge probability as a function of the gain, the measurements have to be performed simultaneously. The pico-amperemeter, used to measure the amplification current (see Section 2.4), registers also discharge events as sharp spikes, altering the gain estimation. To prevent this, an averaging algorithm excludes points that deviate more than from the mean current value. The procedure was used in our previous studies and is explained in more detail in [22]. In each measurement of the discharge probability, up to hundreds of sparks are recorded with the oscilloscope. However, the discharge rate never exceeds to assure reliable gain measurement. The associated statistical error of discharge occurrences is calculated by assuming a Poisson distribution of the latter.
@@ 3.1 Gain comparisons
As expected, the gain dependence on the applied voltage follows an exponential behaviour. A sudden drop of the THGEM gain at higher voltages is a known effect linked to the onset of discharges and discussed in more detail in [22]. When comparing the measurements between different MPGDs and radioactive sources, it can be clearly seen that varying humidity levels in the gas do not significantly affect the absolute gain. Also, no clear dependency can be observed, contrary to the discussion in Section 2.5. However, the measured gain values are not corrected for variations, which, given each measurement is performed on a different day, can dissolve the already vague dependency.
@@ 3.2 Leakage current studies
For the three MPGD structures studied, the leakage current is clearly dependent on gas humidity. Only at very high voltages, beyond the values at which the gain of a structure is measured (i.e., the discharge rate with radioactive sources is too high), is the current/voltage relation non-linear, and the leakage current dependency on the gas humidity is less pronounced. The latter is even reversed in the case of GEM and THGEM, as the highest leakage current values are measured for the dry gas. This can be related to the onset of discharges that are likely to appear at very high fields, especially near the electrode defects. This will be further investigated in the discharge stability studies, discussed in the following Section 3.3.
The study shows that humidity can alter the leakage current measurement, however, in the measured range the differences are marginal. In all cases, the leakage current values measured below the high-voltage/high-current region are at the level of tens of picoamperes. This is well below the limits used in the quality control of MPGDs produced for large-scale projects like ALICE TPC [14], CMS GE1/1 Muon [15] and ATLAS New Small Wheel [16] upgrades during the LHC Long Shutdown 2. Indeed, in all three projects, the HV stability and leakage currents of MPGD structures were measured requiring relative humidity of the gas below 10%, without specifying the exact value. This is well in line with the range considered in this work. As mentioned in Section 1, various studies showed that at RH values above 50% clear instabilities can be observed in the form of enhanced discharge rate or leakage (dark) current values. This region of humidity content is not to be discussed here.
@@ 3.3 Discharge stability studies
Figure 9 shows the discharge probability of the MPGDs using the different sources as a function of the absolute gain. The colour bar indicates the different humidity levels in the gas mixture, while different symbols represent different radioactive sources.
While comparing the results obtained with alpha particles, one can immediately observe that the discharge probability scales inversely with the (TH)GEM hole or MMG cell pitch, thus the hole/cell density. As we have shown in [21, 22, 17] this effect can be explained by simple consideration of a primary charge density arriving at the single amplification cell. Assuming 100% electron collection efficiency (in the case of the MMG1, the gain is corrected for the 100% value - see Section 2.4), the number of primary electrons entering a single amplification hole/cell increases with decreasing hole/cell density. Thus, given the gain value, the probability of reaching the critical charge required for a streamer and spark development is increasing [21, 22, 17]. Measurements with alphas are particularly sensitive to these effects as primary charge densities are very large, and the critical charge limits can be reached already at relatively low gains, where other effects related to, for example, defects, high electric fields at the electrode edges, or contamination, are not yet noticeable.
@@
The right plot in Figure 9 illustrates the discharge rate as a function of the applied amplification voltage for various MPGDs and humidity levels. In addition to the data presented in the left plot of Fig. 9, this plot also includes measurements taken without any radiation source. In this configuration, the observed discharges result from cosmic rays, field emissions, and other factors rather than source irradiation and are referred to as spurious discharges. The data points corresponding to measurements without a source are represented by hollow squares and extend to higher voltages due to reduced primary ionisation. Notably, these measurements reveal a clear correlation between increasing humidity content in the gas mixture and discharge behaviour. Increased stability against spurious discharges is observed in the GEM and THGEM, particularly at the highest voltages. A similar trend is seen in Micromegas, where this dependence is also evident at lower voltages. Additionally, the measurements appear to group at the highest voltages when humidity is introduced, irrespective of its absolute value. This suggests that beyond a certain point, adding more water vapour does not further enhance MPGD stability. These findings indicate that humidity has minimal impact on discharge formation at lower voltages but plays a role in reducing spurious discharges, potentially by mitigating electrode defects or influencing the charging-up of insulating layers.
@@ 4 Conclusion
The impact of humidity on the performance of three different non-resistive MPGD structures (GEM, Thick GEM, and Micromegas) is studied using humidified gas from a dedicated gas system. To evaluate the performance of the detector, different properties are investigated, including gain, energy resolution, leakage current, discharge rate and discharge probability. The measurements are performed both with (alpha, x-ray) and without radiation sources.
Studies on gain and energy resolution indicate that humidity has a negligible impact on detector performance and should not affect investigations on discharge stability. Indeed, the effective Townsend coefficient calculated with Magboltz does not show any dependency on the water content in the considered humidity and electric field ranges. The residual gain variations observed in the 55Fe peak position studies and the absolute gain measurements may be related to the ambient conditions or water absorption in the polyimide material of GEMs. Full interpretation of the effects observed in this work should be considered with a dedicated measurement campaign.
@@
Finally, we observe no major influence of water content on the discharge stability of any of the amplification structures when discharge probability is measured with an alpha or an x-ray source. However, the measurements with 55Fe at the highest amplification fields show a slight dependency towards the higher stability of a humid gas. This observation is further confirmed when the discharge rate is measured without any radiation source. In all the studied structures we observe enhanced stability (lower discharge rates) when the detector is operated with humidified gas. However, no major improvement is observed when the water content is further increased.
We conclude that adding water to the gas mixture increases the discharge stability at the highest fields. The humidity levels do not influence the discharge formation process, however, they reduce the rate of spurious discharges related to electrode defects or charging-up of the insulating layers.

### v2

@@ 1 Introduction
Our measurements are performed in Ar-CO2 (90-10) at the absolute humidity levels of (parts per million, proportion by volume), which correspond to a relative humidity (RH) value of up to 20% at normal temperature and pressure conditions. It should be noted that several groups reported various instabilities in MPGD detectors appearing at much larger relative humidity values while operating/testing a detector structure in an air atmosphere in the laboratory. In [11], instabilities such as increased leakage current and spark rate were observed in GEMs operated in ambient, cleanroom conditions (RH50%). Similar observations have also been made with other types of MPGDs (see, for example, [12] for resistive Micromegas or recent results with µRGroove presented in [13]). For this reason, the humidity limits for the quality control procedures of MPGDs are usually put below the RH value of 50% or even below 10%, as required in the recently concluded large-scale productions for the LHC detector upgrades [14, 15, 16]. This is well in line with the range considered in this work. In the presented studies, we focus on the fundamental stability limits of MPGD-based detectors in a gas mixture with trace amounts of water, and the regions of large RH () are not considered here.
@@ 2 Experimental setup and methods
2.1 Gas mixture
The gas mixture used in the presented studies is Ar-CO2 (90-10). A mixture of argon and carbon dioxide is widely employed in various ratios in modern MPGD-based detectors, including the COMPASS GEM tracker [2], the CMS Muon Endcap GEM [17], and the ATLAS New Small Wheel Micromegas [18], among others. While argon can be considered the most common noble gas in such detectors, the use of CO2 as a quencher — though not always optimal for detector performance — remains highly attractive due to its environmental compatibility (non-greenhouse gas), non-toxicity, non-flammability, and chemical inertness, making it particularly suitable for large-scale MPGD applications. The noble gas to quencher ratio of 90-10 used in this study was primarily chosen to allow direct comparison with a broad range of our previous measurements on MPGD stability (e.g. [19, 20, 21]).
It should be noted that the effects of humidity observed in this study may not be directly transferable to other gas mixtures, as discharge properties and quenching mechanisms can vary significantly. Furthermore, in mixtures containing fluorinated gases, the presence of humidity should be strictly avoided, as it may lead to the formation of hydrofluoric acid, posing a risk to the integrity of detector components [5, 22, 23]. Each gas mixture should therefore be studied individually to assess its specific behaviour under varying humidity conditions.
@@ 2.2 Gas system
A dedicated setup is built to study the effect of humidity on MPGDs, in which a constant amount of water vapour is reliably introduced to the gas. Figure 1 shows a scheme of the humidification process. The original gas mixture flows through a piping system, splitting into two separate lines. One line incorporates a bubbler filled with distilled water, allowing the gas to absorb humidity without introducing additional contaminants like oxygen. Different humidity levels can be achieved by adjusting the flow through the bubbler using a mass flow meter, as shown in Fig. 2. Simultaneously, the second line directs dry gas around the humidifying line until they merge. The gas mixture is continuously flushed through the 4-litre detector vessel at a flow rate of 10 l/h, which is regulated by mass flow controllers and monitored using a rotameter installed in the dry gas line. A Rapidox 3100 Multigas Analyser from Cambridge Sensotec is placed after the detector vessel to record the oxygen and humidity content of the mixture. The manufacturer does not specify the precision of the humidity measurements. However, based on the reproducibility of the values obtained in this work, the uncertainty in the water content is considered negligible. The oxygen content is kept at a minimum ( ppm) throughout all measurements. The ambient conditions are tracked with thermometers and a pressure sensor.
@@
| Pitch | Density | Optical transparency | Producer | |||||||
@@ 2.5 Gain measurements
The absolute gain is defined as the ratio of the amplification current to the primary current measured in the drift volume. The currents are measured using a floating multi-channel picoamperemeter [25]. For the (TH)GEM, this process is carried out with zero induction field, while the amplified signal from avalanche multiplication is read at the bottom of the foil. The primary current is measured at the top (TH)GEM electrode, with the bottom electrode and the anode plane grounded. A drift field of is applied between the cathode and the GEM foil, at which the electron collection efficiency of primary charges is expected to be nearly 100% [19, 20, 26, 27]. The measurement with a Micromegas is performed similarly by applying a drift field of and keeping the mesh and the anode plane grounded. Since the electron transparency of MMG1 is known to be less than 100% for the measured field configurations, the gain values are corrected using data obtained with this specific Micromegas structure [21], where electron transparency is measured as a function of the ratio , with representing the field strength in the amplification region. For MMG3, a 100% collection efficiency can be safely assumed [21].
@@
The similar current values observed under both dry and humid (3000 ppm) conditions, along with the slightly lower current measured at 1500 ppm water content, suggest that the observed variations are not due to a systematically increased humidity level but rather to day-to-day fluctuations in temperature and pressure—especially considering that the differences between the curves remain below 3. All three curves also exhibit similar time constants of approximately 200 seconds. Based on these observations, we conclude that the procedure for voltage application, including the 900-second interval prior to the start of the measurement, can be applied independently of the water content in the gas.
@@ 2.6 Amplitude spectra
The energy resolution values are calculated by fitting a Gaussian to the main peak of the spectrum and taking the ratio of the Gaussian width () to the mean peak position . The systematic uncertainty is estimated by varying the fit ranges. The mean peak position is then corrected for the temperature and pressure () variations, assuming linear dependency of gain (thus value) on ratio, which was monitored throughout the measurement. It should be noted that no detailed calibration is available for the detectors under study. Thus, although the variations were rather limited during the measurement time, the residual gain dependency remains uncorrected.
@@
The GEM gain dependency shown in Fig. 4 is less apparent in the measurement with decreasing humidity. Apart from the residual influence, this may also be related to water absorbed in polyimide during the first measurement series, which influences the electrical and electrostatic properties of the GEM structure (e.g. charging-up behaviour). Indeed, in the case of MMG3, no significant gain dependency and no differences are observed between the measurements with increasing and decreasing humidity content. The Micromegas pillars are made of photopolymer solder mask material (DuPontTM Pyralux® PC 1000) and cover roughly of the detector area, contrary to GEMs, where polyimide foil (Apical®) fills more than 75% of the active area (see Table 1). Thus, the water uptake by the MMG3 detector shall be largely reduced in comparison to GEMs, which may explain the different results obtained with these two types of detectors. In addition, the hygroscopicity of GEM and MMG3 materials shall be considered, as they can further influence the water trapping in these two structures. It is expected that the water absorption of a polymer mask is lower than that of the Apical® foil (water absorption of 2.7%, according to the producer [34]), however, no exact value could be identified for the former.
@@
For completeness, the effect of gas humidity on energy resolution is studied. It is observed to be very weak for both GEM and MMG3, even though, given the electronegativity of water molecules, one could expect the attachment to deteriorate the energy resolution. As discussed above, the influence of humidity on the effective Townsend coefficient for different gases may vary depending on the electric field strength, however, we do not consider this effect in the current studies. The GEM measurements show a slight improvement in the energy resolution with increasing water content, which may be related to the increasing gain values (see above). Still, the relative change of energy resolution by less than 5% in the measured humidity range can be considered marginal. To summarise, the effect of humidity on the performance of the detector is found to be negligible and should not alter the studies on discharge stability. Of course, a complete interpretation of the observed effects should be considered with a dedicated measurement campaign.
@@ 2.7 Discharge measurement
The discharge rate is defined as the number of measured discharges occurring within the measurement time . Consequently, the discharge probability is given by the ratio of the discharge rate to the measured source rate , following:
Discharges are counted by placing an open-ended cable within the detector chamber, where it functions as a simple antenna. The cable is connected to an oscilloscope, which records and displays the discharge signals, as shown in Fig. 6.
To determine the discharge probability as a function of gain, both quantities must be measured simultaneously. The pico-amperemeter, used to measure the amplification current (see Section 2.5), registers also discharge events as sharp spikes, altering the gain estimation. To prevent this, an averaging algorithm excludes points that deviate more than from the mean current value. The procedure was used in our previous studies and is explained in more detail in [20]. In each measurement of the discharge probability, up to hundreds of sparks are recorded with the oscilloscope. However, the discharge rate never exceeds to assure reliable gain measurement. The associated statistical error of discharge occurrences is calculated by assuming a Poisson distribution of the latter.
@@ 3.1 Gain comparisons
As expected, the gain dependence on the applied voltage follows an exponential behaviour. A sudden drop of the THGEM gain at higher voltages is a known effect linked to the onset of discharges and discussed in more detail in [20]. When comparing the measurements between different MPGDs and radioactive sources, it can be clearly seen that varying humidity levels in the gas do not significantly affect the absolute gain. Also, no clear dependency can be observed, contrary to the discussion in Section 2.6. However, the measured gain values are not corrected for variations, which, given that each measurement is performed on a different day, can dissolve the already vague dependency.
@@ 3.2 Leakage current studies
For the three MPGD structures studied, the leakage current is clearly dependent on gas humidity. At very high voltages, specifically those exceeding the operational range used for gain measurements (where the discharge rate with radioactive sources becomes excessively high), the current-to-voltage relationship deviates from linearity. In this regime, the influence of gas humidity on leakage current is significantly reduced. The latter is even reversed in the case of GEM and THGEM, as the highest leakage current values are measured for the dry gas. This can be related to the onset of discharges that are likely to appear at very high fields, especially near the electrode defects. This will be further investigated in the discharge stability studies, discussed in the following Section 3.3.
The study shows that humidity can alter the leakage current measurement, however, in the measured range, the differences are marginal. In all cases, the leakage current values measured below the high-voltage/high-current region are at the level of tens of picoamperes. This is well below the limits used in the quality control of MPGDs produced for large-scale projects like ALICE TPC [14], CMS GE1/1 Muon [15] and ATLAS New Small Wheel [16] upgrades during the LHC Long Shutdown 2. Indeed, in all three projects, the HV stability and leakage currents of MPGD structures were measured, requiring relative humidity of the gas below 10%, without specifying the exact value. This is well in line with the range considered in this work. As mentioned in Section 1, various studies showed that at RH values above 50% clear instabilities can be observed in the form of enhanced discharge rate or leakage (dark) current values. This region of humidity content is not to be discussed here.
@@ 3.3 Discharge stability studies
Figure 9 shows the discharge probability of the MPGDs using the different sources as a function of the absolute gain. The colour bar indicates the different humidity levels in the gas mixture, while different symbols represent different radioactive sources. The large vertical error bars (discharge probability) arise from the limited number of discharge events — typically fewer than ten — recorded during a single 900 s measurement, constrained by the requirement to maintain stable conditions. Regarding the horizontal error bars, in the high-gain regime, the frequent discharging (above 0.3 Hz) makes it difficult to find sufficiently long, stable intervals for reliable current measurements. The limited number of measurement points within these time windows leads to large statistical uncertainties in the gain determination, reflected in the size of error bars.
While comparing the results obtained with alpha particles, one can immediately observe that the discharge probability scales inversely with the (TH)GEM hole or MMG cell pitch, thus the hole/cell density. As we have shown in [19, 20, 21] this effect can be explained by simple consideration of a primary charge density arriving at the single amplification cell. Assuming 100% electron collection efficiency (in the case of the MMG1, the gain is corrected for the 100% value — see Section 2.5), the number of primary electrons entering a single amplification hole/cell increases with decreasing hole/cell density. Thus, given the gain value, the probability of reaching the critical charge required for a streamer and spark development is increasing [19, 20, 21]. Measurements with alphas are particularly sensitive to these effects as primary charge densities are very large, and the critical charge limits can be reached already at relatively low gains, where other effects related to, for example, defects, high electric fields at the electrode edges, or contamination, are not yet noticeable.
@@
Figure 10 illustrates the discharge rate as a function of the applied amplification voltage for various MPGDs and humidity levels. In addition to the data presented in Fig. 9, this plot also includes measurements taken without any radiation source (hollow points). In this configuration, the observed discharges result from cosmic rays, field emissions, and other factors rather than source irradiation and are referred to as spurious discharges. The data points corresponding to measurements without a source are represented by hollow squares and extend to higher voltages due to reduced primary ionisation. Notably, these measurements reveal a clear correlation between increasing humidity content in the gas mixture and discharge behaviour. Increased stability against spurious discharges is observed in the GEM and THGEM, particularly at the highest voltages. A similar trend is seen in Micromegas, where this dependence is also evident at lower voltages. Additionally, the measurements appear to group at the highest voltages when humidity is introduced, irrespective of its absolute value. This suggests that beyond a certain point, adding more water vapour does not further enhance MPGD stability. These findings indicate that humidity has minimal impact on discharge formation at lower voltages but plays a role in reducing spurious discharges, potentially by mitigating electrode defects or influencing the charging-up of insulating layers.
@@ 4 Conclusion
The impact of humidity on the performance of three different non-resistive MPGD structures (GEM, Thick GEM, and Micromegas) is studied using humidified Ar-CO2 (90-10) gas mixture from a dedicated gas system. To evaluate the performance of the detector, different properties are investigated, including gain, energy resolution, leakage current, discharge rate and discharge probability. The measurements are performed both with (alpha, x-ray) and without radiation sources.
Studies on gain and energy resolution indicate that humidity has a negligible impact on detector performance and should not affect investigations on discharge stability. Indeed, the effective Townsend coefficient calculated with Magboltz shows no dependence on the water content within the studied humidity and electric field ranges. However, a complete understanding of this behaviour may require more sophisticated avalanche modelling, including, for example, ion kinetics. The residual gain variations observed in the 55Fe peak position studies and the absolute gain measurements may also be related to the ambient conditions or water absorption in the polyimide material of GEMs. Full interpretation of the effects observed in this work should be considered with a dedicated measurement campaign.
@@
Finally, we observe no major influence of water content on the discharge stability of any of the amplification structures when discharge probability is measured with an alpha or an x-ray source in Ar-CO2 (90-10) mixture. However, the measurements with 55Fe at the highest amplification fields show a slight dependency towards the higher stability of a humid gas. This observation is further confirmed when the discharge rate is measured without any radiation source. In all the studied structures, we observe enhanced stability (lower discharge rates) when the detector is operated with humidified gas. However, no major improvement is observed when the water content is further increased. These findings suggest that the addition of a small amount of water vapour in the range of can effectively enhance discharge stability. Nevertheless, the optimal humidity level should be determined individually for each application and gas mixture, taking into account whether the aim is to suppress leakage currents, mitigate discharge instabilities, or address other specific operational requirements of the detector.
In conclusion, while humidity does not appear to influence the intrinsic discharge formation mechanism, it does reduce the rate of spurious discharges, likely associated with electrode imperfections or the charging-up of insulating surfaces.